Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Probability - Target hit

Expert replies
by nisagl750 » Mon Dec 24, 2012 1:19 am
You have 3 balls and a target to hit. The probability of hitting the target in one try is 90%. What is the probability that the target will be hit in 3 tries.

My approach:
Probability that target will be hit = 1-P(target missed)
P(Target missed) = 1/10 in one try
In three tries = 1/10 * 1/10 * 1/10 = 1/1000

So, probability that the target will be hit = 999/1000

Am I right? Or I am missing something?
Join the discussion
Source: — Problem Solving |

by The Iceman » Mon Dec 24, 2012 2:02 am
You are absolutely right.

It is the same as the probability of getting at least one shot on target, which is 1-(0.1)^3 =0.999=99.9%

Another way to solve this problem is to consider cases in which the target is hit on the very first shot OR in the second shot OR in the third Shot =0.9+0.1*0.9+0.1*0.1*0.9=0.999=99.9%

It is logical to say that since the chance of hitting the target is 90% on one go, the chance increases as the shooter is given more chance to have a go at the target, but it will never attain 100% probability even if infinite arrows were shot. However the chance of success will tend to move very close towards 100% if infinite arrows are shot.
Join the discussion

by nisagl750 » Mon Dec 24, 2012 3:13 am
The Iceman wrote:You are absolutely right.

It is the same as the probability of getting at least one shot on target, which is 1-(0.1)^3 =0.999=99.9%
Thank you.
I was confused because I thought If probability of hitting target in one hit is 90%, how can it be more than 90% in 3 hits.
But after reading your reply I got it. My above doubt stands valid if the question asks the probability of target hit in all three hits. In that case the probability will be well below 90%

9/10 * 9/10 * 9/10 = 729/1000
Join the discussion

by The Iceman » Mon Dec 24, 2012 3:47 am
nisagl750 wrote:
The Iceman wrote:You are absolutely right.

It is the same as the probability of getting at least one shot on target, which is 1-(0.1)^3 =0.999=99.9%
Thank you.
I was confused because I thought If probability of hitting target in one hit is 90%, how can it be more than 90% in 3 hits.
But after reading your reply I got it. My above doubt stands valid if the question asks the probability of target hit in all three hits. In that case the probability will be well below 90%

9/10 * 9/10 * 9/10 = 729/1000
Yes, the odds will go on stacking against the shooter as you have asked too much from him :). If you him to do that in 50 such shots his chance reduces from 90% to 0.52%.
Join the discussion

by Sapana » Fri Dec 28, 2012 9:57 am
It is a good question. Thank you for posting it. but, I do not understand what is difference in calculating 0.9*0.9*0.9 for three hits and calculating it by (1-0.001) [0.001=1/10*1/10*1/10]. I know the answers are different but Can any one please explain the difference in logic?
Join the discussion

by nisagl750 » Sun Dec 30, 2012 10:11 am
Sapana wrote:It is a good question. Thank you for posting it. but, I do not understand what is difference in calculating 0.9*0.9*0.9 for three hits and calculating it by (1-0.001) [0.001=1/10*1/10*1/10]. I know the answers are different but Can any one please explain the difference in logic?
Hi Sapana,

0.9*0.9*0.9 gives the probability that target is hit all three times (i.e all three balls hit the target , Hit Hit Hit)
&
1-0.001 gives the probability that target is hit (i.e target is hit at least once)

I hope you get it. If you have any doubt, please post it.

-
Nisagl750
Join the discussion

by jkaustubh » Sun Dec 30, 2012 11:30 am
The question says that The probability of hitting the target in one try is 90%.

but I see that here we are using the fact that probability of hitting the target is 90%. which may of may not be the same as probability of hitting the target in one try (single shot).

Can somebody please explain this situation!!!
Replying a query takes patience and time. The least a person can do is to thank the reply.
Join the discussion

by nisagl750 » Sun Dec 30, 2012 11:02 pm
jkaustubh wrote:The question says that The probability of hitting the target in one try is 90%.

but I see that here we are using the fact that probability of hitting the target is 90%. which may of may not be the same as probability of hitting the target in one try (single shot).

Can somebody please explain this situation!!!
Hi Kaustubh,

Can you please elaborate a bit on your question. This is what I understand.
If the probability of hitting target in one try is 90% then prob of hitting target in every hit is 90%. It goes on increasing with increase in no. of chances given to the player and goes on decreasing with increase in hit ratio required by the player

i.e If the player is given 10 chances, probability will be much higher than 90% (close to 1)
&
If the player is asked to hit 10 out of 10 chances, the probability is very less.
Did I answer you? Or you have any other doubt which I could not understand?

-
nisagl750
Join the discussion